OpenStudy (anonymous):

5x+10/x+2 * 2x/4x-10 multiply

OpenStudy (anonymous):

\[\frac{ 5x+10 }{ x+2 }*\frac{ 2x }{ 4x-10 }\]

OpenStudy (decentnabeel):

\[\frac{5x+10}{x+2}\cdot \frac{2x}{4x-10}=\frac{5x}{2x-5}\]

OpenStudy (triciaal):